A company produces product X at a unit cost of N10. If fixed costs are N450, 000, and each unit sells for N25,
how many must be sold for the company to make a positive profit of N510,000.

Answer by ikleyn(52776)

.
A company produces product X at a unit cost of N10. If fixed costs are N450, 000, and each unit sells for N25,
how many must be sold for the company to make a positive profit of N510,000.
~~~~~~~~~~~~~~~~


Your equation is


    Revenue - Cost           = Profit,    or

    25x     - (450000 + 10x) = 510000.


Simplify and solve


    25x - 10x = 510000 + 450000

    15x = 960000

     x    = 960000/15 = 64000 units.


ANSWER.  64000 units must be produced and sold.

Solved.


It is very standard problem on profit.